OpenStudy (anonymous):

does this series converge or diverge? sum from n=1 to infinity of log(n*sin(1/n) )

OpenStudy (anonymous):

@hartnn @dan815

OpenStudy (perl):

$$ \Large{\sum_{n=1}^{\infty} \log(n\cdot \sin(1/n) ) }$$

OpenStudy (perl):

By the comparison test, the series converges.

OpenStudy (perl):

do you agree that sin(1/n) < 1/n

OpenStudy (perl):



OpenStudy (perl):

it is true that sin x < x for x > 0 , where x is in radians

OpenStudy (perl):

remember, arclength = radius * theta . Here radius is 1 . theta I labeled as x

OpenStudy (perl):

you can see that geometrically, the arc of length x must be longer than sin x .